OpenStudy (anonymous):

***PLEASE HELP*** Last year, Lien earned $1500 more than her husband. Together they earned $37,500. How much did each of them earn? ***I cant figure out the equation!***

OpenStudy (anonymous):

we have answered the question, go check out the older posts :)

OpenStudy (anonymous):

nvm lol, I'll paste my answer here :) Alright, let's take it one step at a time :). Write your given in one side and RTF (required to find) in the other side and you'll have the following as given: 1) Let's let Lien be (L) 2) Husband be (H) Given: Lien has 1500 more than her husband, do we know how much her husband has? no , so we'll keep him as H , but we know that Lien has : L (has) = 1500 ( more) + (than her husband) H Now together they gained 37500 so : L + H = 37500 These are your equations , now you must link them to find how much did each get so you'll have : L = 1500 + H, substitute L in the other equation and you'll get: L + H = 37500 1500 + H + H= 37500 , now solve for H (husband) and you'll know how much he'll get so: H = 18,000, so the husband has $18,000 Now substitute this in the first equation to know how much Lien got so : L = 1500 + H L = 1500 + 18000 L = 19500 So Lien got , $19500 Do you understand what was going on now? ^_^

OpenStudy (anonymous):

Thanks! I kinda understand. I sorta had the numbers and letters right, but my biggest problem is trying to figure out the actual equation to solve for.

OpenStudy (anonymous):

np ^_^